# Efficient Radar Modulation Recognition via a Noise-Aware Ensemble Neural Network

### v2

@@ Abstract
Electronic warfare support (ES) systems intercept adversary radar signals and estimate various types of signal information, including modulation schemes. The accurate and rapid identification of modulation schemes under conditions of very low signal power remains a significant challenge for ES systems. This paper proposes a recognition model based on a noise-aware ensemble learning (NAEL) framework to efficiently recognize radar modulation schemes in noisy environments. The NAEL framework evaluates the influence of noise on recognition and adaptively selects an appropriate neural network structure, offering significant advantages in terms of computational efficiency and recognition performance. Furthermore, we employ feature extraction blocks to enhance the efficiency of the proposed recognition model. We present the analysis results of the recognition performance of the proposed model based on experimental data. Our recognition model demonstrates superior recognition accuracy with low computational complexity compared to conventional classification models.
@@ I Introduction
Although existing CNN-based models exhibit efficiency in radar modulation recognition, a significant issue with these models is their substantial decrease in performance in low SNR environments. This decrease occurred because they employed structures with limited feature extraction capabilities to reduce the computational load. This issue can be addressed by increasing the computational complexity of the recognition model. However, recognition models applied to ES systems need to estimate signal parameters instantaneously. In high SNR environments, recognition models that enhance feature extraction capabilities with a high computational load can result in unnecessary power consumption and excessively long inference times. Consequently, existing models that use fixed neural network architectures generate unnecessary computations in low-noise environments or exhibit insufficient recognition capabilities in high-noise environments.
This study proposes a novel neural network architecture to address these issues, as depicted in Fig. 2(b). The proposed architecture exploits a noise-aware ensemble learning (NAEL) technique that evaluates the influence of noise on recognition and adaptively adjusts the network structure accordingly. The proposed model initially utilizes a lightweight preliminary recognizer to perform modulation recognition. When the recognition outcomes of the preliminary recognizer are compromised by noise interference, an advanced recognizer capable of extracting deeper features is employed to determine the final modulation recognition outcomes. By employing a conditional decision-making model with this noise-aware approach, unnecessary computation in high SNR environments can be reduced, while recognition accuracy in low SNR environments is enhanced.
@@ II-A Intercepted Radar Signal
where $x(n)$ represents the complex envelopes of the transmitted radar signals, $w(n)$ represents the complex additive white Gaussian noise, $A$ is the amplitude, and $f(n)$ and $\phi(n)$ are the frequency and phase modulation functions, respectively.
@@
TABLE I: Frequency and phase modulation functions for 12 modulation schemes.
@@ II-B Time-frequency Analysis
$\begin{split}\textup{CWD}(&t,f)=\int^{\infty}_{-\infty}\exp(-j\omega\tau)\int^{\infty}_{-\infty}\sqrt{\frac{\sigma}{4\pi\tau^{2}}}\\ &\exp\left(-\frac{\left(\mu-t\right)^{2}}{4\tau^{2}/\sigma}\right)y\left(\mu+\frac{\tau}{2}\right)y^{*}\left(\mu+\frac{\tau}{2}\right)d\mu d\tau\end{split}$ (2)
where $t$ and $f$ are the time and frequency variables, respectively, $\omega=2\pi f$ is the angular frequency, and $\sigma\left(\sigma>0\right)$ is a scaling factor determining the trade-off between cross-term suppression and frequency resolution. In this study, we set $\sigma$ to $1$ for CWD time-frequency analysis.
@@ III Proposed Modulation Recognition Model
This section describes the proposed modulation recognition model exploiting the NAEL framework. Fig. 3 illustrates the overall neural network architecture of our proposed recognition model. The model consists of three neural networks: a preliminary recognition network (PRN), noise-aware network (NAN), and advanced recognition network (ARN). A brief description of each network is provided below.
  • The PRN primarily determines the modulation recognition results from the TFI.
  • The NAN assesses whether the modulation recognition results derived by the PRN are affected by noise.
  • To mitigate potential noise-induced misrecognition, the NAEL framework utilizes the ARN, which excels in precise feature extraction, to make the final decision. This occurs when the NAN determines that recognition results are likely to be significantly influenced by noise.
@@ III-A FE Block
Our proposed recognition neural network employs FE blocks following the structures illustrated in Figs. 4 and 5. The structure depicted in Fig. 4 is used repetitively from the second stage onward, while the structure in Fig. 5 is used only in the first stage. In the first stage, depending on the parameters used in the block design, the model can adjust the number of output channels, perform downsampling, or both simultaneously.
@@
Downsampling operations can be integrated into the first stage of the FE block to enhance computational efficiency and extract key features effectively. If the block parameters are set to include a downsampling operation, the structure illustrated in Fig. 5(a) is used. This structure reduces the spatial dimensions of the output feature map by half and adjusts the number of output channels to $C_{o}$. In contrast, the structures illustrated in Fig. 5(b) is used. Here, the FE block retains the original spatial dimensions of the input feature map while adjusting the number of output channels to $C_{o}$.
@@ III-B Modulation Recognition Neural Networks
This subsection provides a detailed explanation of the structures of PRN and ARN that constitute the proposed recognition model. PRN and ARN utilize FE blocks for modulation recognition. Table III summarizes the input dimensions and parameters used for each block, where $C_{o}$ denotes the number of channels in the block output, $r$ represents the number of repetitions for spatial feature extraction, and $s$ indicates the convolution layer stride in the first stage of the block. For the FE block, the structure shown in Fig. 5 is used in the first stage, and the structure in Fig. 4 is repeated $r-1$ times for subsequent stages. In addition, if $s=2$, the structures shown in Fig. 5(a) is used in the first stage; if $s=1$, the structure in Fig. 5(b) is used. Based on our experimental results, setting the parameters listed in Table III yielded the best modulation recognition performance.
TABLE III: Detailed structure of the recognition networks within the NAEL framework
@@
For classification, transforming a feature map into a one-dimensional feature vector is crucial for effective processing by a classifier based on fully connected layers. To achieve this, we employed a channel-expansion (CE) block after the last FE block. This block consists of a 1$\times$1 convolution layer, followed by BN and ReLU6 activation. The CE block enhances the representation of each channel in $\mathcal{F}$, the output of the last FE block, to capture essential information. Subsequently, a one-dimensional feature vector is produced by averaging the values across channels of the CE block output. This feature vector is then fed into a classifier composed of fully connected layers followed by a softmax layers. The classifier computes class scores based on the feature vector, identifying the modulation scheme with the highest score.
@@ III-C Noise-Aware Neural Network
A key proposal of this study involves using a noise-aware method to determine the impact of noise on recognition results derived by the PRN. To achieve this, the NAN generates a gradient map illustrating the regions of the input TFI on which the PRN focuses to derive modulation recognition results. The method for generating the gradient map is based on computing gradient information from the PRN’s feature map $\mathcal{F}$ as gradient information is crucial for visually highlighting the data influencing the decision [36]. To obtain the gradient map, the process starts by computing the gradient between the $c$-th class score $y^{c}$ (the highest among computed class scores) from the PRN’s classifier output, and the $k$-th channel of the feature map $\mathcal{F}^{k}$ from the output of the PRN’s CE block. This gradient $\partial y^{c}/\partial\mathcal{F}^{k}$ represents the influence of $\mathcal{F}^{k}$ on $y^{c}$. To determine the importance of the $k$-th feature map channel for the $c$-th class score, importance weights for the $k$-th channel $w^{c}(k)$ are computed using global average pooling across the width and height dimensions of $\mathcal{F}^{k}$. This process is summarized as follows:
$w^{c}(k)=\frac{1}{{H\times W}}\sum_{t_{i}=1}^{H}\sum_{t_{j}=1}^{W}\frac{\partial y^{c}}{\partial\mathcal{F}^{k}(t_{i},f_{j})}$ (7)
@@
By performing a weighted sum on $\mathcal{F}^{k}$ with the importance weights and applying the ReLU activation function, a gradient map $\mathcal{G}^{c}$ is produced. This gradient map provides a visual explanation for the recognition basis of the $c$-th modulation scheme as follows:
@@
Fig. 6 shows examples of gradient maps for simulated signals in cases of correct classification and misclassification. Costas and T2 code signals were used in an environment with an SNR of -11 dB. In the proposed recognition model, the feature map $\mathcal{F}$ has dimensions of 8$\times$8, resulting in a gradient map of the same 8$\times$8 size. As shown in Fig. 6(e) and (g), during correct classification, the gradient map exhibits the highest values at indices 4 and 5 in the frequency domain, corresponding to the signal’s center frequency. The ES system was assumed to accurately estimate the center frequency for all simulated signals before performing modulation recognition, with a frequency offset applied to align the center frequency with the middle of the TFI frequency axis. In case of misclassification, as shown in Fig. 6(f) and (h), high values appear at positions other than the center frequency. This occurs because the gradient map highlights noisy areas where the signal does not exist, leading the PRN to focus on noise and make incorrect decisions.
Figure 7: Distribution of the $f_{max}$ for the two cases: (a) correctly classified case, and (b) misclassified case.
@@
Fig. 7 shows the distribution of frequency domain values at the highest points in the gradient maps for both correctly and incorrectly classified cases. For this analysis, 50 simulated signals for each of the 12 modulation schemes were generated in an environment with an SNR of -15 dB. The frequency index $f_{max}$ of the highest point in $\mathcal{G}^{c}(t,f)$ was extracted as follows:
$f_{max}=\textup{arg}\max_{f}\left(\sum_{t=1}\mathcal{G}^{c}(t,f)\right).$ (9)
@@
As seen in Fig. 7(a), which shows the distribution of $f_{max}$ when correctly classified, the highest values are concentrated at frequency indices 4 and 5, corresponding to the center frequency. In contrast, Fig. 7(b), which depicts the distribution when misclassified, shows significantly more outliers compared to the correct classification case. These results indicate that using a gradient map can be highly effective in assessing the reliability of recognition results.
Gradient maps serve as visualizations of the neural network’s decision-making process, allowing for the diagnosis of faults using this explainability. The proposed model uses a gradient map to determine whether current recognition results are significantly affected by noise. Fig. 7 demonstrates the usefulness of the gradient map using $f_{max}$; however, a deep-learning-based classifier that utilizes all elements of the gradient map as input can achieve better noise awareness performance. To this end, we flattened the gradient map into a one-dimensional gradient vector of size 64$\times$1. Subsequently, a classifier consisting of fully connected and softmax layers predicts whether the recognition is affected by noise.
TABLE IV: Modulation Parameters for Dataset Generation
@@
Figure 8: Measurement environment for constructing test datasets.

### v3

# This work has been accepted for publication in the IEEE Internet of Things Journal

@@ Abstract
Electronic warfare support (ES) systems intercept adversary radar signals and estimate various types of signal information, including modulation schemes. The accurate and rapid identification of modulation schemes under conditions of very low signal power remains a significant challenge for ES systems. This paper proposes a recognition model based on a noise-aware ensemble learning (NAEL) framework to efficiently recognize radar modulation schemes in noisy environments. The NAEL framework evaluates the influence of noise on recognition and adaptively selects an appropriate neural network structure, offering significant advantages in terms of computational efficiency and recognition performance. We present the analysis results of the recognition performance of the proposed model based on experimental data. Our recognition model demonstrates superior recognition accuracy with low computational complexity compared to conventional classification models.
@@ I Introduction
More recently, several advanced deep learning models have been introduced to enhance radar modulation recognition performance further. LPI-Net[11] employs a time-frequency analysis-based lightweight feature extraction technique within a deep CNN framework to improve the recognition accuracy of LPI radar waveforms. LWCNN[31] focuses on robustness against various types of noise, leveraging squeeze-and-excitation networks to achieve improved classification performance under noisy conditions. Additionally, the vision transformer-based recognition model[32] incorporates transformer architectures instead of conventional CNNs, allowing for better learning of long-range dependencies in modulation characteristics.
Although existing CNN-based models exhibit efficiency in radar modulation recognition, a significant issue with these models is their substantial decrease in performance in low SNR environments. For instance, MobileNet, which utilizes depth-wise convolutions, and ShuffleNet, which employs channel shuffling, both adopt lightweight architectures that prioritize computational efficiency; however, they lack the flexibility to adapt the depth of their feature extraction when noise levels increase. This results in a significant drop in recognition accuracy under low-SNR conditions, as their shallow feature representations struggle to capture crucial modulation characteristics. This issue can be addressed by increasing the computational complexity of the recognition model. However, recognition models applied to ES systems need to estimate signal parameters instantaneously. In high SNR environments, recognition models that enhance feature extraction capabilities with a high computational load can result in unnecessary power consumption and excessively long inference times. Consequently, existing models that use fixed neural network architectures generate unnecessary computations in low-noise environments or exhibit insufficient recognition capabilities in high-noise environments.
This study proposes a novel neural network architecture to address these issues, as depicted in Fig. 2(b). The proposed architecture exploits a noise-aware ensemble learning (NAEL) technique that evaluates the influence of noise on recognition and adaptively adjusts the network structure accordingly. The proposed model initially utilizes a lightweight preliminary recognizer to perform modulation recognition. When the recognition outcomes of the preliminary recognizer are identified as affected by noise interference, an advanced recognizer capable of extracting deeper features is employed to determine the final modulation recognition outcomes. By employing a conditional decision-making model with this noise-aware approach, unnecessary computation in high SNR environments can be reduced, while recognition accuracy in low SNR environments is enhanced. Compared to existing models, which maintain a fixed lightweight architecture, the proposed model adapts its feature extraction depth dynamically, ensuring optimal trade-offs between computational efficiency and recognition accuracy. This adaptability makes it particularly effective for real-time ES applications, where both performance and efficiency are critical.
@@ II-A Intercepted Radar Signal
where $n$ is the discrete-time index, $x(n)$ represents the complex envelopes of the transmitted radar signals, $w(n)$ represents the complex additive white Gaussian noise, $A$ is the amplitude, and $f(n)$ and $\phi(n)$ are the frequency and phase modulation functions, respectively.
@@ II-B Time-frequency Analysis
$\begin{split}\textup{CWD}(&t,f)=\int^{\infty}_{-\infty}\exp(-j\omega\tau)\int^{\infty}_{-\infty}\sqrt{\frac{\sigma}{4\pi\tau^{2}}}\\ &\exp\left(-\frac{\left(\mu-t\right)^{2}}{4\tau^{2}/\sigma}\right)y\left(\mu+\frac{\tau}{2}\right)y^{*}\left(\mu-\frac{\tau}{2}\right)d\mu d\tau\end{split}$ (2)
where $t$ and $f$ are the time and frequency variables, respectively, $\omega=2\pi f$ is the angular frequency, $\sigma\left(\sigma>0\right)$ is a scaling factor determining the trade-off between cross-term suppression and frequency resolution. In (2), the parameter $\mu$ represents the temporal reference point at which the signal characteristics are analyzed, while $\tau$ denotes the time lag used to evaluate the local correlation properties of the signal around $\mu$.
@@ III Proposed Modulation Recognition Model
This section describes the proposed modulation recognition model exploiting the noise-aware ensemble learning (NAEL) framework. This framework consists of three neural networks: the preliminary recognition network (PRN), noise-aware network (NAN), and advanced recognition network (ARN). Their respective roles are summarized in Table III. The interaction and decision-making flow among the PRN, NAN, and ARN are illustrated clearly in Fig. 3. Furthermore, Fig. 4 presents the detailed neural network architecture of our proposed recognition model, explicitly showing the specific structure and connections of the PRN, NAN, and ARN.
TABLE III: Description of networks and their specific roles within NAEL
Figure 3: A flowchart illustrating the interactions between PRN, NAN, and ARN in the NAEL framework.
@@ III-A FE Block
Our proposed recognition neural network employs FE blocks following the structures illustrated in Figs. 5 and 6. The structure depicted in Fig. 5 is used repetitively from the second stage onward, while the structure in Fig. 6 is used only in the first stage. In the first stage, depending on the parameters used in the FE block design, the model can adjust the number of output channels, perform downsampling, or both simultaneously.
@@
Downsampling operations can be integrated into the first stage of the FE block to enhance computational efficiency and extract key features effectively. If the block parameters are set to include a downsampling operation, the structure illustrated in Fig. 6(a) is used. This structure reduces the spatial dimensions of the output feature map by half and adjusts the number of output channels to $C_{o}$. Otherwise, the structure illustrated in Fig. 6(b) is used. Here, the FE block retains the original spatial dimensions of the input feature map while adjusting the number of output channels to $C_{o}$.
This FE block structure is particularly advantageous for radar modulation recognition due to its ability to efficiently extract time-frequency features while maintaining computational efficiency. By employing depth-wise convolution, the model effectively captures spatial dependencies within individual feature map channels. The subsequent point-wise convolution integrates the extracted features across channels, enabling the model to learn complex radar modulation patterns more effectively. Additionally, the incorporation of skip connections helps preserve critical modulation characteristics, ensuring robust recognition performance even in low-SNR conditions. With these design choices, the proposed framework can effectively identify modulation schemes while balancing accuracy, efficiency, and robustness.
@@ III-B Modulation Recognition Neural Networks
This subsection provides a detailed explanation of the structures of PRN and ARN that constitute the proposed recognition model. PRN and ARN utilize FE blocks for modulation recognition. Table IV summarizes the input dimensions and parameters used for each block, where $C_{o}$ denotes the number of channels in the block output, $\alpha$ implies channel expansion factor, $r$ represents the number of repetitions for spatial feature extraction, and $s$ indicates the convolution layer stride in the first stage of the block. Based on our experimental results, setting the parameters listed in Table IV yielded the best modulation recognition performance. For the FE block, the structure shown in Fig. 6 is used in the first stage, and the structure in Fig. 5 is repeated $r-1$ times for subsequent stages. In addition, if $s=2$, the structures shown in Fig. 6(a) is used in the first stage; if $s=1$, the structure in Fig. 6(b) is used.
TABLE IV: Detailed structure of the recognition networks within the NAEL framework
@@
For classification, transforming a feature map into a one-dimensional feature vector is crucial for effective processing by a classifier based on fully connected layers. To achieve this, we employed a channel expansion (CE) block after the last FE block. This block consists of a 1$\times$1 convolution layer, followed by BN and ReLU6 activation. The CE block enhances the representation of each channel in $\mathcal{F}$, the output of the last FE block, to capture essential information. Subsequently, a one-dimensional feature vector is produced by averaging the values across channels of the CE block output. This feature vector is then fed into a classifier composed of fully connected layers followed by a softmax layer. The classifier computes class scores based on the feature vector, identifying the modulation scheme with the highest score.
@@ III-C Noise-Aware Neural Network
A key proposal of this study involves using a noise-aware method to determine the impact of noise on recognition results derived by the PRN. To achieve this, the NAN generates a gradient map illustrating the regions of the input TFI on which the PRN focuses to derive modulation recognition results. The method for generating the gradient map is based on computing gradient information from the PRN’s feature map $\mathcal{F}$, as gradient information is crucial for visually highlighting the data influencing the decision [38]. To obtain the gradient map, the process starts by computing the gradient between the $c$-th class score $y^{c}$ (the highest among computed class scores) from the PRN’s classifier output, and the $k$-th channel of the feature map $\mathcal{F}^{k}$ from the output of the PRN’s CE block. This gradient $\partial y^{c}/\partial\mathcal{F}^{k}$ represents the influence of $\mathcal{F}^{k}$ on $y^{c}$. To determine the importance of the $k$-th feature map channel for the $c$-th class score, an importance weight $w^{c}(k)$ is introduced. This weight quantifies how significantly the $k$-th channel contributes to recognizing the $c$-th modulation class and is computed using global average pooling across the width and height dimensions of $\mathcal{F}^{k}$. This process is summarized as follows:
$w^{c}(k)=\frac{1}{{H\times W}}\sum_{j=1}^{H}\sum_{i=1}^{W}\frac{\partial y^{c}}{\partial\mathcal{F}^{k}(t_{i},f_{j})}$ (7)
@@
By performing a weighted sum on $\mathcal{F}^{k}$ with the importance weights and applying the ReLU activation function, a gradient map $\mathcal{G}^{c}$ is produced. The gradient map $\mathcal{G}^{c}$ visually highlights the time-frequency regions in the input TFI that significantly contribute to the PRN’s recognition of the $c$-th modulation scheme. The gradient map is computed as follows:
@@
Gradient maps can serve as visualizations of the neural network’s decision-making process, allowing for the diagnosis of faults using this explainability [39]. To explicitly evaluate the influence of noise on recognition, the proposed method analyzes the gradient map $\mathcal{G}^{c}$. Specifically, our proposed method exploits the property that wireless signals occupy regions near the center frequency when analyzed in the frequency domain. Therefore, if high gradient values appear predominantly in noise-dominant regions rather than around the signal’s center frequency, it indicates that the recognition process can be significantly influenced by noise. This relationship arises because the distribution of the gradient map is directly influenced by the SNR. As the SNR increases, the network can focus more accurately on the signal components, resulting in a gradient map with high energy concentrated at the actual signal locations (i.e., regions around the center frequency). Conversely, when the SNR decreases, noise can be misinterpreted as signal features, leading to energy spreading in the gradient map beyond the true signal regions. This phenomenon introduces ambiguity in the recognition process, making it crucial for the NAN to assess the reliability of the recognition results based on gradient map characteristics.
@@
Fig. 7 shows examples of gradient maps for simulated signals in cases of correct classification and misclassification. Costas and T2 code signals were used in an environment with an SNR of $-11$ dB. In the proposed recognition model, the feature map $\mathcal{F}$ has dimensions of 8$\times$8, resulting in a gradient map of the same 8$\times$8 size. As shown in Fig. 7(e) and (g), during correct classification, the gradient map exhibits the highest values at indices 4 and 5 in the frequency domain. The ES system was assumed to accurately estimate the center frequency for all simulated signals before performing modulation recognition, with a frequency offset applied to align the center frequency with the middle of the TFI frequency axis. Therefore, for correctly classified cases, the highest gradient values are concentrated at indices 4 and 5, which correspond to the signal’s center frequency. In the case of misclassification, as shown in Fig. 7(f) and (h), high values appear at positions other than the center frequency. This occurs because the gradient map highlights noisy areas where the signal does not exist, leading the PRN to focus on noise and make incorrect decisions.
@@
Fig. 8 shows the distribution of $f_{max}$ for both correctly and incorrectly classified cases. For this analysis, 50 simulated signals for each of the 12 modulation schemes were generated in an environment with an SNR of $-15$ dB. The frequency index $f_{max}$ represents the frequency location where the gradient map $\mathcal{G}^{c}(t,f)$ has the highest activation, indicating the frequency region most influential for modulation recognition. The $f_{max}$ can be obtained as follows:
$f_{max}=\textup{arg}\max_{f}\left(\sum_{t=1}^{W}\mathcal{G}^{c}(t,f)\right).$ (9)
@@
As seen in Fig. 8(a), which shows the distribution of $f_{max}$ when correctly classified, the highest values are concentrated at frequency indices 4 and 5, corresponding to the center frequency. In contrast, Fig. 8(b) illustrates the distribution in cases of misclassification, revealing a noticeably higher number of outliers compared to when classifications are correct. This difference highlights a clear distinction between correct and incorrect recognition results. These findings strongly indicate that utilizing a gradient map is highly effective in evaluating the reliability of recognition outcomes, as it helps detect anomalies associated with misclassification.
The NAN framework quantifies noise impact by analyzing the distribution of activation in the gradient map. Specifically, when the PRN correctly classifies a signal, the highest activation values in the gradient map $\mathcal{G}^{c}(t,f)$ tend to be concentrated at the center frequency, indicating that the model focuses on meaningful signal features. However, in cases of misclassification, these activations are more dispersed, suggesting that the model is influenced by noise rather than the actual signal. To systematically assess this effect, we measure the spatial distribution of high-activation areas within the gradient map and compare them with expected signal characteristics.
Figure 8: Distribution of the frequency index $f_{max}$ in (a) correct and (b) incorrect classifications at an SNR of $-15$ dB.
Fig. 8 demonstrates the usefulness of the gradient map using $f_{max}$; however, a deep learning-based classifier that utilizes all elements of the gradient map as input can achieve better noise awareness performance. While using $f_{max}$ as a decision criterion provides a straightforward approach to assessing noise influence, it only considers the frequency index with the highest activation in the gradient map. This approach does not fully capture the spatial distribution of activations across the entire time-frequency representation. In contrast, by utilizing the entire gradient map as input, the proposed classifier can learn more complex patterns associated with noise influence, leading to improved classification performance in diverse noise environments. Therefore, the NAN framework further refines noise assessment by using a deep learning-based classifier. Instead of relying solely on a single frequency index, the classifier takes the entire gradient map as input, allowing it to learn more complex patterns of noise interference.
The gradient map is first flattened into a one-dimensional feature vector with 64 elements. It is then processed through two fully connected layers, which sequentially expand the feature dimension to 256 and then to 512. Each layer applies batch normalization and a ReLU6 activation to improve learning stability. Finally, a fully connected layer maps the 512-dimensional feature vector to two output neurons, representing the classification outcomes: reliable (not noise-affected) and unreliable (noise-affected). A softmax activation function is applied to the output layer to convert the final classification scores into probability values, ensuring that the model assigns a confidence score to each possible outcome. If the classifier assigns a higher probability to the reliable category, the PRN’s recognition result is accepted as the final decision. If the unreliable category receives a higher probability, ARN is activated to conduct a more advanced feature extraction process and refine the recognition result.
The use of fully connected layers ensures that the classifier does not rely solely on a single frequency index but instead learns to extract discriminative features from the entire gradient map. Additionally, the softmax classifier assigns probabilistic confidence scores to the predictions, allowing the system to dynamically adapt to varying SNR conditions. Unlike fixed-threshold methods, which may not generalize well across different noise environments, the learned model effectively captures non-linear dependencies, resulting in a more robust noise-aware classification. This classification process inherently accounts for the impact of varying SNR levels as the network learns to recognize the structural differences in gradient maps across different SNR environments. By analyzing how the gradient map’s energy distribution shifts under different SNR conditions, the NAN can dynamically adapt its noise-aware decision-making, leading to improved recognition robustness even in challenging scenarios.
TABLE V: Modulation Parameters for Dataset Generation